What mass of glucose, C&H12O6, must be dissolved in 150.0 g of water so that the mole fraction of glucose is 0.125?
Answer #1

Mole fraction of glucose = 0.125

Then mole fraction of water = 1 - 0.125 = 0.875

Molar mass of glucose = 180 g/mole

Calculate the moles of water as follows:

Mwater= mass/molar mass = 150/18 moles.

Calculate the mass of glucose as follows:

Xwater = Mwater/(Mwater+Mglucose)

0.875 = (150/18)/((150/18)+(x/180))

x = 214.286 gram.

Therefore, mass of glucose required is 214.286 grams.

Please give good rating.